OpenStudy (anonymous):

need to write this complex number in to rectangular form. 6(cos3pi/2+i sin3pi/2)

jimthompson5910 (jim_thompson5910):

that number is in the form \[\Large r*( \cos(\theta) + i*\sin(\theta) )\]

jimthompson5910 (jim_thompson5910):

to convert to rectangular or cartesian form, you need to find x and y the good news is that you can use these formulas \[\Large x = r*\cos(\theta)\] \[\Large y = r*\sin(\theta)\]

OpenStudy (perl):

all you need to do is evaluate your expression

OpenStudy (perl):

what is cos (3pi/2)

jimthompson5910 (jim_thompson5910):

oh true, you could also express it in a+bi form

OpenStudy (perl):

cos(3pi/2) = 0 sin(3pi/2) = -1 6( 0 + i*-1) = 0 -6i

OpenStudy (anonymous):

thank you guys so much, Jim i found your example on algebra.com that helped me out tremendously

jimthompson5910 (jim_thompson5910):

I'm glad it helped

OpenStudy (perl):

what example on algebra.com?